OpenStudy (anonymous):

36 is 72% of what number?

OpenStudy (anonymous):



OpenStudy (anonymous):

....how did you get that?

OpenStudy (anonymous):

is means =

OpenStudy (anonymous):

of means times

OpenStudy (anonymous):

number means x

OpenStudy (anonymous):

solve for x

OpenStudy (akashdeepdeb):

@wilks311 That would give 36 is 72 parts of which number. Here we have to calculate the percentage.

OpenStudy (akashdeepdeb):

@Bunbun1234321 Do you know percentages?

OpenStudy (anonymous):

Ok.... And yes

OpenStudy (anonymous):

So would you do is over over = percent over 100?

OpenStudy (anonymous):

Oops I meant is over of

OpenStudy (akashdeepdeb):

Yes Here it asks us 36 is 72 percent of what number. So there is a number greater than 36, and it forms 72% of that number. :)

OpenStudy (akashdeepdeb):

Do you know what is 50 % of 36?

OpenStudy (anonymous):

Yes it's 3.6 right?

OpenStudy (anonymous):

Oops nevermind

OpenStudy (anonymous):

It's 18

OpenStudy (anonymous):

Sorry I gtg thanks for your help!

OpenStudy (akashdeepdeb):

Not quite. No problem. :) See, the very concept of maybe a bit counterintuitive at first but it is quintessential that you get it. So here goes... When I asked you for 50% of 36 I essentially asked you to divide 36 into hundred parts and take only 50 out of the 100 parts of 36. So when you divide 36 into hundred parts you get 36/100 = 0.36 So we have to take 50 of these '0.36's. 50 * 3.6 = 18 And that is how we get 18.

OpenStudy (akashdeepdeb):

Coming to your question, it asks us 36 is 72% OF WHAT NUMBER. So let that number be 'num' [Just for ease in understanding] So the question if converted to a mathematical statement would be... 72 percent of 'num' = 36 So, if num is divided into 100 parts, 72 of those parts give 36. So if num is divided into 100 parts it'll give ==> num/100 [This is just a way to understand it, don't do it in an exam. ;-) ] \[\frac{72 * num}{100} = 36\] By cross multiplication in equations \[num = \frac{36*100}{72}\] So \(num = 50\) That just tells you that from a bigger number, 36 forms 72% of it and thus 50 just forms 100% of that number, which is the number itself. :D Did you understand this? If you didn't get it, I don't mind explaining once more. :)